# Automatic Intermodal Loading Unit Identification using Computer Vision: A Scoping Review

### v1

@@ 4.1 Methods
Figure 1: Approach categories used for ILU identification task over the years.
@@
In the subsequent five years, a modest increase in the use of ML can be observed, particularly in hybrid DIP and ML approaches [106, 107, 108, 109, 110, 12]. This reflects the growing recognition of ML’s potential for enhancing robustness in character and symbol recognition tasks [111] in CV. However, the field lagged several years behind the broad adoption of DL for other applications, despite its transformative influence on CV [112, 113] from 2012 onwards (e.g., the emergence of AlexNet [114]). This delay may be attributed to the scarcity of labelled ILU identification datasets, limited computational infrastructure, and the conservative approach for adopting new technologies in the logistics industry.
@@
ILU identification with fixed cameras often resembles OCR from scanned documents, as the background is mostly uniform. Segmenting and recognising characters is a sensible approach in such cases. Until 2016, all methods developed for ILU identification focused on character-level localisation and recognition, and this approach is still in use with DL methods [122, 123, 124]. However, with the increasing use of mobile cameras (e.g. vehicle-mounted cameras), the task has evolved to resemble scene-text detection and recognition. As a result of this, the trend is shifting towards using comprehensive scene-text detection and recognition methods rather than relying on character detection and recognition.
By 2021–2025, DL-based approaches had established themselves as the predominant paradigm for ILU identification. Even hybrid DIP-based methods virtually disappeared, indicating that the DL models now offer both the required accuracy and operational robustness [66, 125, 126]. The relative infrequency of hybrid approaches suggests that most effective solutions now rarely require DIP or traditional ML methods, except in specialised scenarios such as real-time constraints [104, 16, 18, 127, 124, 15].
4.2 Main Approaches for Deep Learning
Observing the growing interest of DL methods, we now focus on the CV tasks relevant for extracting textual information. OCR techniques are designed to identify and localize individual characters in an image. They are used predominantly for document text recognition and perform well under standard conditions, such as uniform backgrounds and consistent lighting (e.g., scanned documents). However, these techniques may fall short when dealing with text-in-wild scenarios [128], where the background and lighting are highly diverse and the text orientation is arbitrary.
For tasks involving the diverse image properties of natural scenes, such as identifying signs in urban areas [129], two primary subtasks are distinguished. Scene-text detection focuses on identifying and localizing text regions within an image without interpreting their content, while scene-text recognition converts these localized regions into machine-readable characters [130, 131, 132]. The combination of these tasks is known as scene-text spotting, which handles text detection and recognition in an end-to-end manner. These methods are generally faster and more efficient due to their unified feature extraction module [133, 134].
Focusing solely on character detection and recognition can be insufficient when image conditions are highly diverse, such as varying lighting and backgrounds. Relying on character detection in diverse scenes can increase false positives, as many patterns may look similar to letters. Additionally, DL methods require vast amounts of data, and character-level annotation is more expensive than word-level or line-level annotation due to its granularity.
Table 2: Main approaches to ILU identification. The table reports how often each combination appears in the surveyed literature. $\bullet$ = detection, $\bullet$ = recognition.
@@ 4.2 Main Approaches for Deep Learning
20 54.1%
4 10.8%
3 8.1%
2 5.4%
2 5.4%
2 5.4%
1 2.7%
1 2.7%
1 2.7%
1 2.7%
This shift from character detection and recognition to text detection and recognition is shown in Table 2, where the majority of articles employ DL methods for both scene-text detection and recognition tasks. The share of hybrid approaches using non-DL methods for one of these steps are minimal. Examples include using traditional techniques for text detection such as cropping a predefined area for searching ID codes [127] or for character recognition, template matching after a DL model detects text areas [135]. Instead, researchers largely favour more comprehensive pipelines that rely on DL in each step. This shift towards unified DL-based approaches for both scene-text detection and recognition highlights their potential. However, to further enhance efficiency, researchers should focus more on text spotting [136, 134].
In ILU identification, text recognition remains the greatest challenge. Environmental conditions and technical issues such as unreliable hardware or connectivity problems between devices can cause delays at the gate. These delays increase waiting times for stakeholders, lead to congestion, add to environmental pollution, and create further setbacks for subsequent loading units. When recognition errors occur, the risk of incorrect inventory lists rises, which not only requires time-consuming manual corrections in the handling process but can also raise security concerns [14]. To address these challenges more effectively, end-to-end approaches have been proposed. By combining detection and recognition into a single forward pass with a unified feature extraction module, they streamline the process and improve overall efficiency.
Few studies introduce object detection as a prior step to scene-text detection and recognition [73, 80]. Detecting the ILU first reduces the search area for the text detection step and enables association of detected and recognised IDs with other objects, thereby improving scene understanding [73].
@@ 4.3 Datasets
Datasets form the backbone of reliable ILU detection. The quality, diversity, and annotation of the underlying data strongly influence how well models perform once deployed in real-world port environments. Carefully curated datasets with consistent labelling not only enable robust training and evaluation, but also determine whether methods can generalise beyond laboratory settings.
Understanding the properties of datasets is therefore essential for evaluating the performance and applicability of CV-based ILU identification systems. Below, we summarise the main characteristics and settings of the datasets used in the eligible studies, highlighting camera positioning, diversity, and availability.
4.3.1 Image Acquisition Setups
@@
In contrast, mobile cameras are mounted on unmanned aerial vehicles, reach stackers, terminal trucks, or other industrial vehicles. For ground-level perspectives, we also considered handheld images of ILUs, as they closely resemble those captured by vehicle-mounted systems. Unlike fixed setups, mobile cameras provide flexibility by moving across the terminal and capturing a wider range of perspectives. This mobility enables real-time port monitoring and localisation of each ILU in real-world coordinates, as well as monitoring their positions by handling vehicles such as reach stackers [138, 139]. While they cover wider spatial areas and provide richer scene diversity, mobile cameras introduce more complex image variability, requiring advanced processing methods and diverse publicly available datasets.
@@
Camera position Only Fixed 47 74.60 %
Fixed and Ground 10 15.87 %
Fixed, Aerial and Ground 3 4.76 %
Only Ground 1 1.58 %
Only Aerial 1 1.58 %
Not reported 1 1.58 %
Availability Not public 54 85.71 %
Fully available 3 4.76 %
Partially available 3 4.76 %
Upon request 3 4.76 %
Behind paywall 1 1.58 %
As shown in Table 3 the most common collection setting is the fixed CCTV cameras at OCR-gates [140, 123, 141] or cranes (74.60 %) [80, 108] and a mixture of these images with mobile phone snapshots (15.87 %) [67]. However, vehicle-mounted cameras are increasingly used alongside fixed cameras to capture wider areas and provide information across the terminal [142, 84, 122, 73].
@@ 4.3.2 Diversity
Most datasets consist of images gathered from the real-world settings without controlled environments such as the World’s busiest ports like Shanghai [16], Singapore [44], Hong Kong [143], Kaohsiung [13], and Hamburg [73], as well as smaller inland ports like Dresden and Riesa [84]. Besides the terminal-specific data collection, web scraping were also used [125, 72]. In addition, controlled environments such as the Shanghai Maritime University test facility [42] were used for data collection.
Other datasets were created with images taken in Malaysia [66], Sweden [123], Vietnam [18, 124, 127, 68], Thailand [20], South Korea [142, 71, 144, 108, 102], Japan [99], and Italy [145]. However, datasets consisting of images from multiple locations are rare. For DL methods, capturing the nuances of ILUs across different countries and their respective ports can be important.
The number of images used in the experiments of the eligible articles ranges from a minimum of 30 images to a maximum of 34000, with a median of 1050 images. Since the DIP-based methods does not require a huge amount of data, researchers who tackled this problem before DL, experimented with a lower number of images compared to recent studies. However, with the emergence of ML and DL, the need for data has increased significantly. Most of the studies have addressed this need by pre-training their models on general-purpose scene-text detection and recognition datasets such as ICDAR2015 [146], CTW1500 [147], and SynthText [148], then fine-tuning on smaller ILU datasets.
Early datasets already included efforts to represent diverse conditions in order to evaluate generalisation. Lighting variability is the most discussed factor, with datasets often including images taken in bright, low-light, and uneven lighting scenarios [100, 43]. Considering how the earlier, thresholding-based methods may get affected by small changes in light intensity, diversifying the lighting conditions in the experiments was a major focus [145, 97, 143, 96]. However, currently lighting conditions are classified with more general classes: Day and night settings are frequently included to test the adaptability of the systems [104, 122, 73, 16].
Weather conditions such as rain, snow, and fog are also represented, as they can significantly affect image quality and ILU identification performance [73, 67, 126, 143]. Additionally, some data sets feature images with motion blur, camera noise, dirty, rusty, damaged, or partially occluded ID codes to further challenge algorithms with the exceptional conditions that exist in real-world settings [104, 73, 105, 126, 72]. Rarely were these conditions simulated to increase the data sets [125].
@@ 4.3.3 Availability
Table 3 shows the availability of the datasets used by the identified articles. The significant majority of the datasets are not publicly available, which creates a substantial barrier for research and development in this area. The lack of availability also prevents the reproducibility of results and fair comparison of the proposed methods. Further details for the publicly available datasets are shown in Table 4.
Among these, the TRansportation Unit Detection and Identification (TRUDI) dataset [73] stands out for its annotation density. It offers 35,034 instances across 733 images, covering five classes: Container, trailer, tank container, ID code, and logo. The Shipping Container Code dataset [23] leads in the number of images, and the Container Number-OCR dataset [105], though smaller (2,851 images from Tianjin Port), offers similar challenging conditions as others: variable weather conditions and degraded ID codes (dirt, rust, damage). Uniquely, TRUDI introduces motion blur as an additional challenging condition, as most of its images were captured via UAVs, reach stackers, and terminal trucks.
@@
Shipping container code [23] 3600 4000 Day/night, Rainy, snowy Dirty, damaged ID Not reported
Container Number-OCR [105] 2851 2851 Low/high light Foggy Dirty, damaged ID Tianjin Port
TRUDI [73] 733 35034 Day/night Rainy, snowy, foggy Motion blur Dirty, damaged ID
@@ 4.4 Evaluation
To evaluate the proposed ILU identification methods, several evaluation metrics are commonly used. As there are multiple stages of ILU identification using CV methods, such as text detection and text recognition, each stage can be evaluated individually. However, end-to-end accuracy or success rate, is the most relevant metric due to its practical importance. It is the most strict metric as one misrecognised character is enough for failed identification.
Since the ILU identification problem is mostly solved with multiple stages such as text detection and text recognition, some articles only report the results of individual stages and do not include end-to-end results. However, the results for individual stages offer limited insight for the applicability of the method for the ILU identification task.
@@
For object detection or text detection tasks, mean average precision (mAP) and average precision (AP) are commonly used, measuring precision across various intersection over union (IoU) levels. The IoU metric is important for assessing the overlap between predicted and actual bounding boxes. As for the recognition, besides precision, recall and F1-score, other metrics such as the character error rate (CER) is used. CER measures the proportion of characters that were incorrectly predicted compared to the total number of characters in the reference text, using the Levenshtein-distance [149].
@@
Although the recent studies use the same, established metrics adapted from general purpose text detection and recognition task, putting the reported results for ILU identification into a frame is not possible with the limited amount of publicly available datasets. Due to different datasets used in the experiments, the reported end-to-end accuracy of the proposed methods ranges from 5 % [142] to 96,17 % [144]. This level of variance can be observed in every metric, including the processing speed.
@@ 4.5 Future Research Directions
In the literature, future directions for the ILU identification task mainly focus on enhancing the datasets by including more images taken from diverse environments such as various distances, angles, and under different weather and lighting conditions. This expansion aims to cover multiple ports and scenarios to improve model generalisation. However, increasing the diversity of proprietary datasets would not solve the fundamental problem with the datasets for ILU identification. Recent papers acknowledge the lack of benchmark datasets and its negative effects on the field [122, 73]. Given that creating a new dataset for downstream CV tasks are tedious, time consuming, and costly [150, 151], alternatives such as synthetic data generation can be considered [152, 148], e.g., by using a game engine [153].
Handling challenging conditions, such as low-light scenarios, occlusions, and damaged ID codes, is another direction pointed by the researchers, with techniques proposed to manage partial occlusions and degraded text [105]. Current solutions are relying on ”off-the-shelf” super-resolution methods such as Super-Resolution-CNN [154]. Literature on scene-text super-resolution has focused primarily on natural language, leaving an opportunity for research specifically on contextless text such as ILU codes. Alternatively, multi-view recognition is also identified as a future direction due to its potential against degraded ID codes [23, 12] as multiple images of a single ILU are captured from different angles in this approach.
Improving image quality through preprocessing is another focus area. Suggested methods include image enhancement techniques such as noise reduction [17], or the use of Generative Adversarial Networks [155, 156] to address issues such as shadows, rust, and dirt [23, 105]. Postprocessing steps are also recommended to verify recognition results by using the check digit and the owner code. [15, 157, 69, 72].
For DL-based methods, refining network architectures to increase the receptive field so that the model can see more of the scene in a single pass. It is important for detecting containers appears to be small due to perspective [122]. Another future work suggestion is incorporating attention mechanisms to boost recognition accuracy [158]. The development of unified, end-to-end models is suggested to streamline processing pipelines [73, 140, 105, 157].
@@
Improvements for real-time processing are also highlighted, with a focus on optimising algorithms for mobile devices and enhancing systems to work in real-time, potentially through video stream analysis [123, 42].

### v2

@@ 4.1 Methods
Figure 4: The number of publications and their approach categories used for ILU identification task over the years.
In the subsequent five years, a modest increase in ML-based methods can be observed, particularly in hybrid DIP and ML approaches [108, 109, 110, 111, 112, 8]. This reflects the growing recognition of ML’s potential for enhancing robustness in character and symbol recognition tasks in CV [113]. However, the field lagged several years behind the broad adoption of DL, despite its transformative influence on CV solutions [114, 115] from 2012 onwards (e.g., the emergence of AlexNet [116]). This delay may be attributed to the scarcity of labelled ILU identification datasets, limited computational infrastructure, lack of technology skills in the workforce, and the conservative approach for adopting new technologies in the logistics industry [117, 118].
@@
Table 2: Stages defined in DL-based pipelines for ILU identification. The table reports how often each combination appears in the surveyed literature. $\bullet$ = detection, $\bullet$ = recognition.
@@
20 54,1%
4 10,8%
3 8,1%
2 5,4%
2 5,4%
2 5,4%
1 2,7%
1 2,7%
1 2,7%
1 2,7%
Until 2016, all methods developed for ILU identification focused on character-level localisation and recognition with fixed camera settings. This approach is still in use with DL methods [16, 125, 126] as ILU identification on fixed cameras often resembles OCR on scanned documents. However, with the increasing use of vehicle-mounted cameras, the task has evolved to resemble scene-text detection and recognition rather than OCR. As a result of this, the trend is shifting towards using comprehensive scene-text detection and recognition methods rather than relying on character detection and recognition. Table 2 shows that the majority of articles employ DL methods for both scene-text detection and recognition tasks. The share of hybrid approaches using non-DL methods for one of these steps are minimal. Examples include using traditional techniques for text detection such as cropping a predefined area for searching ID codes [127] or for character recognition, template matching after a DL model detects text areas [128]. Instead, researchers largely favour more comprehensive pipelines that rely on DL in each step.
By 2021–2025, DL-based approaches had established themselves as the predominant paradigm for ILU identification. Even hybrid DIP-based methods virtually disappeared, indicating that the DL models now offer both the required accuracy and operational robustness [73, 129, 130]. The decline of hybrid approaches suggests that most effective solutions now rarely require DIP or traditional ML methods, except in specialised scenarios such as real-time constraints [106, 13, 19, 127, 126, 12]. Researchers also address real-time constraints by focusing on text spotting as these methods are generally faster and more efficient due to their unified feature extraction module [131, 132, 133]. A few studies introduce object detection or instance segmentation as a prior step to text detection and recognition [17, 84]. Detecting the ILU reduces the search area for the text detection step and enables association of detected and recognised IDs with other objects, thereby improving scene understanding [17].
@@ 4.2 Datasets
The quality and diversity of annotated data strongly influence how well DL models perform once deployed in real-world port environments. Carefully curated datasets with consistent labelling not only enable robust training and evaluation, but also determine whether methods can generalise beyond laboratory settings. Understanding the properties of datasets is therefore essential for evaluating the performance and applicability of CV-based ILU identification systems. Below, we summarise the main characteristics of the datasets used in the eligible studies, highlighting camera position settings, diversity, and availability.
4.2.1 Image Acquisition Settings
@@
In contrast, mobile cameras are mounted on unmanned aerial vehicles, reach stackers, terminal trucks, or other industrial vehicles and smart devices. For ground-level perspectives, we also considered handheld cameras, as they closely resemble those captured by vehicle-mounted systems. Unlike fixed setups, mobile cameras provide flexibility by moving across the terminal and capturing a wider range of perspectives. This mobility introduce higher image variability, requiring advanced processing methods and diverse datasets. When these challenges addressed, there is a potential for more precise and real-time terminal monitoring. Thanks to sensor equipped vehicles [135, 83], CV pipelines beyond ILU identification can be created. For example, localisation of ILUs through pose estimation, autonomous path optimisation for vehicles navigating around terminals, multi-agent coordination (collaboration of aerial and ground vehicles) for terminal monitoring.
@@
Camera position Only Fixed 47 74,60 %
Fixed and Ground 10 15,87 %
Fixed, Aerial and Ground 3 4,76 %
Only Ground 1 1,58 %
Only Aerial 1 1,58 %
Not reported 1 1,58 %
Availability Not public 54 85,71 %
Fully available 3 4,76 %
Upon request 3 4,76 %
Partially available 2 3,17 %
Behind paywall 1 1,58 %
As shown in Table 3 the most common collection setting is the fixed CCTV cameras at OCR-gates [136, 125, 137] or cranes (74,60 %) [84, 110] and a mixture of these images with smartphone images (15,87 %) [74]. However, vehicle-mounted cameras are increasingly used alongside fixed cameras to capture wider areas and provide information across the terminal [14, 15, 16, 17].
@@ 4.2.2 Diversity
Most datasets consist of images gathered from ports, reflecting their operational settings. Images in these datasets were taken from the world’s busiest ports like Shanghai [13], Singapore [42], Hong Kong [138], Kaohsiung [9], and Hamburg [17], as well as other ports in Malaysia [73], Sweden [125], Vietnam [19, 126, 127, 75], Thailand [21], South Korea [14, 78, 139, 110, 104], Japan [101], and Italy [140]. While rare, there are datasets with images collected in smaller inland ports like Dresden and Riesa [15, 17]. Besides the terminal-specific data collection, web scraping were also used [129, 79]. In addition, controlled environments such as the Shanghai Maritime University test facility [40] were used for data collection. However, datasets consisting of images from multiple locations are rare. For DL methods, capturing the nuances of ILUs across different countries and their respective ports is important.
The number of images used in the experiments in the eligible articles ranges from a minimum of 30 images to a maximum of 34.000, with a median of 1.050 images. Since the DIP-based methods do not require a large amount of data, researchers who tackled ILU identification task before DL, experimented with a lower number of images compared to recent studies. However, with the emergence of ML and DL, the need for data has increased significantly. Most of the studies have addressed this need by pre-training their models on general-purpose scene-text detection and recognition datasets such as ICDAR2015 [141], CTW1500 [142], and SynthText [143], then fine-tuning on smaller ILU datasets [16, 17].
Starting from the early datasets, efforts to represent diverse conditions have been made in order to evaluate generalisation. Lighting variability is the most discussed factor, with datasets often including images taken in bright, low-light, and uneven lighting scenarios [102, 41]. Considering how the earlier, thresholding-based methods get affected by small changes in light intensity, diversifying the lighting conditions in the experiments was a major focus [140, 99, 138, 98]. However, currently lighting conditions are classified with more general classes: Day and night settings are frequently included to test the adaptability of the systems [106, 16, 17, 13].
Weather conditions such as rain, snow, and fog are also represented, as they can significantly affect image quality and ILU identification performance [17, 74, 130, 138]. Additionally, some datasets feature images with motion blur, camera noise, dirty, rusty, damaged, or partially occluded ID codes to further challenge algorithms with the conditions that exist in real-world settings [106, 17, 107, 130, 79]. Rarely were these conditions simulated to increase the diversity of datasets [129].
@@ 4.2.3 Availability
We divided datasets into 5 different categories: Fully available [24, 107, 17], upon request [144, 16, 136], partially available [145, 79], behind paywall [14] and not public. Table 3 shows the availability of the datasets used by the identified articles. The significant majority of the datasets are not publicly available, which creates a substantial barrier for research and development in this area. We were not able to confirm the availability of the datasets that claim to be available upon request as our requests went unanswered. Recent papers acknowledge the lack of benchmark datasets and its negative effects on the field [16, 17]. The lack of availability prevents the reproducibility of results and fair comparison between the proposed methods for researchers. As for the practitioners, while large seaports typically have extensive propriety datasets, publicly available datasets extend these capabilities to smaller terminals, facilitating data-driven optimisation.
After screening the literature, we could find only 3 publicly available datasets. Among these, the TRansportation Unit Detection and Identification (TRUDI) dataset [17] stands out for its annotation density. It offers 35,034 instances across 733 images, covering five classes: Container, trailer, tank container, ID code, and logo. It introduces motion blur as an additional challenging condition, as most of its images were captured via UAVs, reach stackers, and terminal trucks. The Shipping Container Code dataset [24] leads in the number of images, and the Container Number-OCR dataset [107], though smaller (2.851 images from Tianjin Port), offers similar challenging conditions as others: variable weather conditions and degraded ID codes (dirt, rust, damage). Further details for the publicly available datasets are shown in Table 4.
Adopting these publicly available datasets for benchmarking will be very helpful for both researchers and practitioners to evaluate and compare proposed methods.
@@
Shipping container code [24]
3600
4000
Day/night, Rainy, snowy Dirty, damaged ID Not reported
Container Number-OCR [107]
2851
2851
Low/high light Foggy Dirty, damaged ID Tianjin Port
TRUDI [17]
733
35034
Day/night Rainy, snowy, foggy Motion blur Dirty, damaged ID
@@ 4.3 Evaluation
To evaluate the proposed ILU identification methods, several evaluation metrics are commonly used. As there are multiple stages of ILU identification using CV methods, such as text detection and text recognition, each stage can be evaluated individually. In the literature, 60.18 % articles reported the results of individual stages only and did not include end-to-end results. However, the results for individual stages offer limited insight for the applicability of the method for the ILU identification task. End-to-end accuracy or success rate is the most relevant metric due to its practical importance. It is the most strict metric as one misrecognised character is enough for failed identification.
@@
For object detection, instance segmentation or text detection tasks, average precision (AP) and mean average precision (mAP) are commonly used, measuring precision across various intersection over union (IoU) levels. The IoU metric is important for assessing the overlap between predicted and ground truth bounding boxes. As for the recognition, besides precision, recall and F1-score, other metrics such as the character error rate (CER) or word accuracy are used. CER measures the proportion of characters that were incorrectly predicted compared to the total number of characters in the reference text, using the Levenshtein-distance [146].
@@
Although the recent studies use the same, established metrics adapted from general purpose text detection and recognition task, putting the reported results for ILU identification into a frame is not possible with the lack of an widely adopted benchmark dataset and the lack of end-to-end results. Only around 40 % of the studies reported end-to-end accuracy and it ranges from 5 % [14] to 99,51 % [74], with the mean of 84.19 % and standard deviation of 24.02 percentage points. Considering the high variation and the low number of reported end-to-end results ($n=25$), we did not perform further statistical analysis as these will not yield in statistically significant results.
@@ 4.4 Future Research Directions
In the literature, future directions for the ILU identification task mainly focus on enhancing the datasets by including more images taken from diverse environments, introducing different challenges such as various distances, angles, weather and lighting conditions. This expansion aims to cover multiple ports and scenarios to improve model generalisation. However, increasing the diversity of proprietary datasets would not solve the fundamental problem with the datasets for ILU identification. Given that creating a new dataset for downstream CV tasks are tedious, time consuming, and costly [147, 148], alternatives such as synthetic data generation can be considered [149, 143], e.g., by using a game engine [150].
Handling challenging conditions, such as low-light scenarios, occlusions, and degraded or damaged ID codes, is another direction pointed by the researchers. Suggested methods include image enhancement techniques such as noise reduction [18], or the use of Generative Adversarial Networks [151, 152] to address issues such as shadows, rust, and dirt [24, 107]. Super-resolution [153] is a task that can tackle the problems with degraded or damaged ID codes. However, the literature on scene-text super-resolution has focused primarily on natural language, leaving an opportunity for research specifically on contextless text such as ILU codes. Alternatively, multi-view ILU identification is another future direction due to its potential against degraded ID codes [24, 8] as multiple images of a single ILU are captured from different angles in this approach. Postprocessing steps are also recommended to verify recognition results by using the check digit and the owner code. [12, 154, 76, 79].
For DL-based methods, refining network architectures to increase the receptive field so that the model can see more of the scene, especially smaller objects in a single pass [16]. Another future work suggestion is incorporating attention mechanisms to boost recognition accuracy [155] or optimising the tokenisation process of vision transformer-based architectures as there are always the same amount of characters expected for ISO6346-compliant ID codes. For better applicability, the development of unified, end-to-end models is suggested to streamline processing pipelines [17, 136, 107, 154]. While DL models achieve high accuracy, their computational demands often limit deployment on edge devices like drones. Future work should focus on model compression, quantisation, or neural architecture search to enable real-time ILU identification in resource constrained environments.